Solve the equation.
x/3 - x/9 = 8

Answer by lexee25(1)

You need to get the greatest common factor to get rid of the denomenator. multiply x/3 by 9 x/9 by 3 and the 8 you will have to multiply it by 9 and 3.
(9)x/3-(3)x/9=8(3)(9)
then you get 9x-3x=216
add like terms 6x=216
divide 6 on both sides 6x/6=218/6
x=36
Check:
36/3-36/9=8
12-4=8
8=8